OpenStudy (anonymous):

sqrt(8)-sqrt(3)^2 = anyone?

OpenStudy (maheshmeghwal9):

What does this mean? the same question again!

OpenStudy (maheshmeghwal9):

k!np:) it's ur matter:) best of luck:)

OpenStudy (lgbasallote):

why did you repost this?

OpenStudy (anonymous):

to see if anyone had a difference response

OpenStudy (anonymous):

Thanks alot though@ mahesmeghwal9

OpenStudy (lgbasallote):

is it written as \[\sqrt 8 - \sqrt 3^2\] or \[(\sqrt 8 - \sqrt 3)^2\]

OpenStudy (anonymous):

like the bottom one

OpenStudy (lgbasallote):

ugh..i hate it when it'slike that...you are familiar with the formula \[(a - b)^2 = a^2 - 2ab + b^2?\]

OpenStudy (anonymous):

yes i have seen it before

OpenStudy (lgbasallote):

therefore \[(\sqrt 8 - \sqrt 3)^2 = (\sqrt 8)^2 - 2(\sqrt 8)(\sqrt 3) + (\sqrt 3)^2\]

OpenStudy (lgbasallote):

now do you see why i hate it =_=

OpenStudy (lgbasallote):

anyway tell me what you got from simplifying that

OpenStudy (anonymous):

yup and i get so confused too lol

OpenStudy (anonymous):

well that the thing i dont know how to work these problems out

OpenStudy (lgbasallote):

hint! \[(\sqrt a)^2 = a\] \[\sqrt b \times \sqrt c = \sqrt{bc}\]

OpenStudy (lgbasallote):

apply those in the problem

OpenStudy (maheshmeghwal9):

sorry I got wrong catching over the question. @lgbasallote is very intelligent:) really appreciate it:)

OpenStudy (anonymous):

thanks to both of you alls help ill continue figuring out cause i got to go

OpenStudy (lgbasallote):

bye and goodluck :D